OpenStudy (anonymous):

solve 1/5(5x-1)=3 which answer is right 1. x=2.80 2. x=4.00 3. x=-2.80 4. x=3.20

OpenStudy (dangerousjesse):

Thank you

OpenStudy (phi):

can you multiply both sides by 5? (that means write 5* next to each side) what do you get ?

OpenStudy (dangerousjesse):

Multiply both sides by a constant to simplify the equation. Multiply both sides of \[\frac{(5 x-1)}{5} = 3 ~by~ 5:\] \[\frac{5 (5 x-1)}{5} = 5×3\]

OpenStudy (dangerousjesse):

Cancel the common terms in the numerator and denominator after you've done that. You should get\[\frac{5(5x-1)}{5}=\frac{5}{5}*(5x-1)=5x-1: \]\[5x-1=5*3\]

OpenStudy (dangerousjesse):

Then you need to multiply five and three together. What's your equation now?